OpenStudy (anonymous):

-3(2w - 1 ) = 5w - 19 Solve and show work

OpenStudy (agreene):

-3(2w-1)=5w-19 distribute the -3 (-3*2)w +3 = 5w - 19 -6w + 3 = 5w - 19 add 6w and 19 to both sides 22 = (5+6)w 22 = 11w divide both sides by 11 22/11 = w 2 = w